Firm break even when?
1
Expert's answer
2020-12-31T09:41:03-0500

ANSWER

The breakeven point is the output level at which total revenue (TR) is equal to the total expenses (TC).

Hence, we can compute this point using cost types (fixed and variable).

The total revenue (denoted by TR) can be calculated by means of the following formula:

"TR(q)=P*q"

The total costs (denoted by TC) can be calculated as follows:

"TC(q)=FC+VC(q)=FC+AVC*q"

where FC os the fixed cost, AVC - average variable costs ("AVC=VC(q)\/q" ).

Accordind to the breakevenpoint definition, the breakevenpoint (BEP) is the q that is satisfies the following equation:

"TR(BEP)=TC(BEP)"

Hence,

"P*BEP=AVC*BEP+FC"

Then,

"BEP*(P-AVC)=FC"

Hence,

"BEP=FC\/(P-AVC)" .